News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 19 April 2002 9.00am EST

Last week the observing schedule was interrupted due to high radiation associated with solar flare activity. The loads were halted on April 17 at 8:55am EDT through a ground command to execute the SI Safing SCS 107. This ensured that the accumulated radiation dose for ACIS remained below the allowed threshold. Loads were rebuilt and uplinked late on March 17, when the radiation environment had reached an acceptable level. The replan included an initial maneuver to avoid an Earth blockage. 39.2 ks of the observation of Field-142549+353248 was impacted by the event.

Of note last week was the upgrade to a new version of the GRETA ground system software and operating system. The new version (1.2.1) supports the new OS and includes a number of enhancements and bug fixes. In addition, a new version of the Off Line System (9.2B) has been approved for operations that will significantly improve the momentum management process during scheduling.

April 18 marked Chandra's 1000th day in orbit!

The observing schedule for the next week is shown below and includes an observation of GX 1+4 coordinated with RXTE.

------------------------------------------
RCS1325+2858            ACIS-S       Apr 22
HS 1229+4807            ACIS-S
Radiation Belts                      Apr 23
NGC2362                 ACIS-S/HETG
1145-676                ACIS-S       Apr 24
1046-409                ACIS-S
Vela Remnant            HRC-I
FTM1012+2825            ACIS-S
SDSS075652+450258       ACIS-S       Apr 24
PSS0808+5215            ACIS-S       Apr 25
3C 265                  ACIS-S
Radiation Belts
Speedy Mic              ACIS-S/HETG  Apr 26
GX 1+4                  ACIS-S/HETG
Speedy Mic              ACIS-S/HETG  Apr 27
1739-304                ACIS-S
GRS 1758-258            HRC-I
HS 2251+2941            ACIS-S
PKS2201+044             ACIS-S
1655+077                ACIS-S       Apr 28

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports